OpenStudy (anonymous):

[7,8] [2] [9 ] help please with this matrix

OpenStudy (anonymous):

are you trying to multiply the matrices?

OpenStudy (anonymous):

yes bt how would i multiply them?

OpenStudy (anonymous):

ooooo ok thnk you

OpenStudy (anonymous):

ooops sorry. my bad. that is wrong

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

http://www.mathwarehouse.com/algebra/matrix/multiply-matrix.php there are examples on there.

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

1x2 matrix multiplied by a 2x1 matrix gives a 1x1 matrix. so the answer is 7 x 2 + 8 x 9 = 14+72 = [86]

OpenStudy (anonymous):

so how would i do it if i was multiplying a 2x3 matrix to a 3x2 matrix

OpenStudy (anonymous):

you would get a 2x2 matrix. see that link. press next button in each problem. they show the steps.

OpenStudy (anonymous):

ok